zarkam21 (zarkam21):

If the sum of interior angle measures of a polygon is 900 degrees, how many sides does the polygon have

OpenStudy (shelbywyatt):

@InstagramModel

OpenStudy (instagrammodel):

For a polygon with x interior degrees the number of sides 'n' is found by the formula (n-2)180 = x We have (n-2)180 = 900 n-2 = 900/180 n-2 = 5 n = 7 The polygon has 7 sides

OpenStudy (shelbywyatt):

thank you

OpenStudy (instagrammodel):

You're welcome (: